OpenStudy (anonymous):

The slope of the line below is 2. Use the coordinates of the labeled point to find the point-slope equation of the line. (-2,-5) those are the points i was given A. y + 5 = -2(x + 2) B. y - 5 = 2(x - 2) C. y - 5 = -2(x - 2) D. y + 5 = 2(x + 2) someone please help!

OpenStudy (daniellelovee):

this is point slope formula y1-y=m(x1-x) y and x are the points and m is still the slope

OpenStudy (anonymous):

Would it be D?

OpenStudy (daniellelovee):

you get too confused with the formulas hold up let me rewrite it in a different way

OpenStudy (anonymous):

ok.. Sorry i am not good at math.

OpenStudy (daniellelovee):

does it make more sense now?

OpenStudy (anonymous):

kind of so it would be B?

OpenStudy (anonymous):

Could you run through another one like this? but walk me step by step just to make sure i understand it correctly.?

OpenStudy (daniellelovee):

actually nvm lol you were right the first time and I was wrong

OpenStudy (daniellelovee):

it is D because minus and a negative become a plus

OpenStudy (daniellelovee):

and sure

OpenStudy (anonymous):

so its y+5=2(x+2) correct?

OpenStudy (daniellelovee):

yes

OpenStudy (anonymous):

Sweet!!! 6 more questions on this test and its done xD

OpenStudy (daniellelovee):

ok lol you better get an A on that test

OpenStudy (anonymous):

i won't... I got 5 wrong.. So I got everything else fine. but if i get 1 more wrong i don't think i'll pass and i dont want to do this test again.

OpenStudy (daniellelovee):

my God are those 5 questions part of the one I helped you with?

OpenStudy (anonymous):

Did you help me with one about what are the correct statements bout slope formula?

OpenStudy (daniellelovee):

I think so (I wasnt sure about C)

OpenStudy (anonymous):

C was wrong.. It was Rise/run and y1-y0/x1-x0 those were the only 2 answers i needed. Otherwise everything else was fine!

OpenStudy (anonymous):

but it doesn't give me credit if i miss one. it marks the whole problem wrong...

OpenStudy (anonymous):

other than that you have been a great help!

OpenStudy (daniellelovee):

wow ok I feel guilty now xD but either way ask your teacher and give her the same reasons why I think C should have been part of it and maybe you either prove him/her wrong or get at least credit

OpenStudy (anonymous):

Ok, i'll try to reason with the internet. I take it online. lol dont feel guilty. xD

OpenStudy (daniellelovee):

oh virtual classes... well still try to reason with whoever gave you the lesson on slope before you took the test. And post the next question lol lets try to get a passing grade at least

OpenStudy (anonymous):

ok hold on let me get the next question! lol

OpenStudy (anonymous):

this is my only virtual class though